# Localized Schrödinger Bridge Sampler

### v1

@@ 2. Schrödinger bridge sampler
While (14) works well for low-dimensional problems, applications to medium- and high-dimensional problems have remained an open challenge since accurate approximations of the Schrödinger bridge problem require an increasing number of samples as the dimension $d$ of the sample space $\mathbb{R}^{d}$ increases (cf. [21]).
@@
The key observation of this paper is that the approximation of conditional expectations (4) via Schrödinger bridges does not necessarily require an accurate approximation of the Markov chain (8) and that localization can be applied provided conditional independence can be established. This idea will be developed in the following section.
@@ 3. Localized Schrödinger bridge sampler
To introduce the main idea of localizing the Schrödinger bridge sampler developed in [5] we first consider an illustrative example of sampling from a multivariate Gaussian distribution. We will see that localization allows to significantly reduce the number of samples required to achieve a certain accuracy. In particular, the number of samples required to achieve a certain accuracy does not depend on the intrinsic dimension of the samples but rather on the correlation length which typically is much smaller.
@@ 3.1. Motivational example: Gaussian setting
Let $\Delta_{h}\in\mathbb{R}^{d\times d}$ denote the standard discrete Laplacian over a periodic domain $[0,1]$ with mesh-size $h=1/d$. We assume that the sampling distribution $\pi(x)$ is Gaussian with zero mean and covariance matrix
(15) $\Sigma=(I-\Delta_{h})^{-1}.$
Instead of the distribution $\pi(x)$, we are given $M$ samples $x^{(j)}\sim{\rm N}(0,\Sigma)$, $j=1,\ldots,M$, and denote their $\alpha$-th entry by $x_{\alpha}^{(j)}$ for $\alpha=1,\ldots,d$. The goal is to produce more samples from ${\rm N}(0,\Sigma)$ using the time-stepping scheme (14) without making explicit reference to the unknown covariance matrix $\Sigma$. This particular setting of a generative model can become arbitrarily challenging for decreasing mesh-sizes $h=1/d$.
@@
(16) $X(n+1)=X(n)-\epsilon(I-\Delta_{h})X(n)+\sqrt{2}\,\Xi(n),\qquad\Xi(n)\sim{\rm N}(0,\epsilon I).$
@@
and periodic extension of $X_{\alpha}$ for $\alpha=0$ and $\alpha=d+1$. We assume that the step size $\epsilon$ is chosen such that $w_{0}\geq 0$. The EM update (17) shows the conditional expectation value of $X_{\alpha}(n+1)$ only depends on the value of the neighboring grid points of $X(n)$ with weights $w_{0}$ and $w_{\pm 1}$; i.e.,
@@
$d_{\alpha}=3$, to denote the set of neighboring grid points of $X_{\alpha}$.
To help the reader navigating the various indices and sub- and superscripts we summarize here our notation. Superscripts $(j)$ are reserved to denote samples $j=1,\ldots,M$ as well as components of vectors in $\mathbb{R}^{M}$. For example, the components of the probability vector $w\in\mathbb{R}^{M}$ are denoted by $w^{(j)}$. The Greek subscript $\alpha$ with $\alpha=1,\ldots,d$ is reserved to denote components of a vector $x$ in state space $\mathbb{R}^{d}$, i.e. $y_{\alpha}$ for $\alpha=1,\ldots,d$. Subscripts ${[\alpha]}$ are reserved to denote localisation around a component $\alpha$; i.e., $x_{[\alpha]}\in\mathbb{R}^{d_{\alpha}}$.
@@
for $\alpha=1,\ldots,d$. We know from our previous considerations that the conditional expectation value of $X_{\alpha}(n+1)$ should depend on $X_{{[\alpha]}}(n+1/2)$ only. Hence the question arises if we can find appropriately localized probability vectors $w(x)$ for the Schrödinger bridge sampler (14)? The following formal argument can be made. We restrict ${\rm N}(0,\Sigma)$ to $\mathbb{R}^{d_{\alpha}}$ and truncate the samples $x^{(j)}$, $j=1,\ldots,M$, accordingly to yield $x_{{[\alpha]}}^{(j)}$. The covariance matrix $\Sigma_{\rm r}\in\mathbb{R}^{d_{\alpha}\times d_{\alpha}}$ of the reduced random variables $X_{{[\alpha]}}\in\mathbb{R}^{d_{\alpha}}$ is simply the restriction of $\Sigma$ to the corresponding sub-space, which in this particular example is independent of $\alpha$. Furthermore, using the Schur complement, one finds
(22) $\Sigma_{\rm r}^{-1}=\left(\begin{array}[]{ccc}\ast&-\frac{1}{2h^{2}}&\ast\\ -\frac{1}{2h^{2}}&1+\frac{1}{h^{2}}&-\frac{1}{2h^{2}}\\ \ast&-\frac{1}{2h^{2}}&\ast\end{array}\right),$
where $\ast$ denotes entries which differ from the matrix which would be obtained by restricting $\Sigma^{-1}$ to the corresponding sub-space. The important point is that the central elements remain identical (cf. (18)) and that only those entries enter the approximation of the conditional expectation value (19).
@@
(23) $(T_{\alpha})_{jk}=\exp\left(-\frac{1}{4\epsilon}\|x_{[\alpha]}^{(j)}-x_{[\alpha]}^{(k)}\|^{2},\right),\qquad i,j=1,\ldots,M,$
for fixed $\alpha\in\{1,\ldots,d\}$. For each of these localized matrices $T_{\alpha}$ we employ the Sinkhorn algorithm to obtain the Sinkhorn weights $v_{\alpha}\in\mathbb{R}^{M}$ for $\alpha=1,\ldots,d$. The key point is that the Euclidean inner product in $\mathbb{R}^{d}$, $d\gg 1$, is replaced by an inner product in $\mathbb{R}^{d_{\alpha}}$ with $d_{\alpha}=3$. Furthermore, in this particular example, the corresponding Schrödinger bridge couples the restricted Gaussian distribution ${\rm N}(0,\Sigma_{r})$ with itself. Next the single $M$-dimensional probability vector (11) is replaced by $d$ $M$-dimensional probability vectors
@@
(26) $w^{(j)}_{\alpha}(x^{(i)}_{[\alpha]})=(P_{\alpha})_{ij},\qquad i,j=1,\ldots,M,$
@@
We finally obtain the localized approximation
(27) $m_{\alpha}(x_{[\alpha]})=\mathcal{X}_{\alpha}\,w_{\alpha}(x_{[\alpha]}),\qquad\alpha=1,\ldots,d,$
@@
(28) $\mathcal{X}_{\alpha}=(x_{\alpha}^{(1)},\ldots,x_{\alpha}^{(M)})\in\mathbb{R}^{1\times M},$
and the update (14b) is replaced by the localized update
(29) $X_{\alpha}(n+1)=m_{\alpha}(X_{[\alpha]}(n+1/2)),\qquad\alpha=1,\ldots,d.$
In other words, we have replaced a single Schrödinger bridge update in $\mathbb{R}^{d}$ by $d$ Schrödinger bridge updates in $\mathbb{R}^{d_{\alpha}}$. We also introduce the localized data matrix
@@ 3.1.1. Numerical illustration
(31) $x^{(j)}\sim{\rm N}(0,\Sigma)$
for $j=1,\ldots,M$ with a $d\times d$ covariance matrix of the form (15) with a tridiagonal precision matrix with $\Sigma^{-1}_{i,i}=2$, $\Sigma^{-1}_{i,i\pm 1}=-0.5$, and periodic conditions $\Sigma^{-1}_{1,d}=\Sigma^{-1}_{d,1}=-0.5$. The corresponding entries of the covariance matrix are $\Sigma_{i,i}\approx 5.8$, $\Sigma_{i,i\pm 1}\approx 1.6$, $\Sigma_{i,i\pm 2}\approx 0.04$, and $\Sigma_{i,i\pm 3}\approx 0.01$.
We employ the localized Schrödinger bridge sampler with a localisation set comprised of four neighbouring grid points, i.e. $d_{\alpha}=5$. We further employ a standard Schrödinger bridge sampler without localisation. In Figure 1 we compare the generated new samples with the given samples for both sampling strategies. We show the resulting empirical histograms as well as the rows of the empirical covariance matrix $\hat{\Sigma}$. The rows are centered at the middle point using periodicity. To achieve reasonable accuracy for the standard sampler we require $M=5,000$ samples and a bandwidth of $\epsilon=0.25$ for $d=9$. The localised sampler instead achieves superior samples with only a tenth of the data with $M=500$ and $\epsilon=0.025$. Moreover, the accuracy does not depend on the dimension for the localized Schrödinger bridge sampler. We show an example with dimension $d=101$ which allows for the same accuracy as the smaller $d=9$-dimensional case. The agreement with the associated entries $\Sigma_{i,i+j}$, $i=51$, $j\in\{0,\pm 1,\pm 2,\pm 3\}$ of the covariance matrix $\Sigma$ is remarkable. We remark that $d=\mathcal{O}(10^{2})$ is out of range for the standard unlocalised sampler.
Figure 1. Comparison of the samples obtained from the standard and from the localized Schrödinger bridge sampler and given samples drawn from a $d$-dimensional multivariate Gaussian with a localized tridiagonal precision matrix. We show the centred rows of the empirical covariance matrix $\hat{\Sigma}$ (top row) and empirical histograms obtained from using all $d$ components (bottom row). The blue markers denote the empirical covariance for the given samples; the magenta markers show the average over all $d$ rows. For the $d=9$ dimensional multivariate Gaussian we show results for the standard sampler trained on $M=5,000$ samples (left column) and the localised sampler trained on $M=500$ samples (middle column). The localised sampler was also run for a $d=101$-dimensional multivariate Gaussian with only $M=500$ samples (right column).
@@ 3.2. Localized Schrödinger bridge sampler for general measures
The strategy of constructing a dimension-reduced localised Schrödinger bridge sampler as presented in the previous example of a multivariate Gaussian readily extends to general priors.
@@
The key assumption we make is that of conditional independence of $x_{\alpha}$ on $x_{\backslash{[\alpha]}}$, which allows for the dimension reduction. The conditional expectation value (27) turns out to be a Monte-Carlo approximation of the conditional expectation under this assumption. More precisely, given the transition density of overdamped Langevin dynamics, denoted here by $p(x^{\prime}|x;\epsilon)$, we obtain
@@
where the second line follows from the conditional independence assumption. Here $p_{\alpha}(x_{\alpha}|x;\epsilon)$ denotes the conditional distribution in $x_{\alpha}$. It is reasonable to assume that we can construct a reversible overdamped Langevin process with invariant distribution $\pi_{\alpha}(x_{[\alpha]})$ and transition kernel $p_{\alpha}(x_{[\alpha]}^{\prime}|x_{[\alpha]};\epsilon)$ on $\mathbb{R}^{d_{\alpha}}$. Here $\pi_{\alpha}(x_{[\alpha]})$ denotes the marginal distribution of $\pi(x)$ in $x_{[\alpha]}$. Then detailed balance of this dimension-reduced Langevin process is given by
@@
Note that in our localized Schrödinger bridge sampler detailed balance is ensured by the Sinkhorn algorithm which renders the Markov chain reversible. Detailed balance then implies for the conditional expectation value
@@
We note that $\rho_{\alpha}(x_{[\alpha]}^{\prime}|x_{[\alpha]};\epsilon)$ is a density with respect to the reference measure induced by $\pi_{\alpha}(x_{[\alpha]})$. We finally approximate the integral in (34c) via Monte Carlo approximation using the restricted data samples $x_{[\alpha]}^{(j)}\sim\pi_{\alpha}$, $j=1,\ldots,M$; i.e.,
@@
with $w_{\alpha}^{(j)}(x_{[\alpha]})\propto\rho_{\alpha}(x_{[\alpha]}^{(j)}|x_{[\alpha]};\epsilon)$ such that $\sum_{j}w_{\alpha}^{(j)}(x)=1$. This expression is of the form used in the localized Schrödinger bridge sampler (27). However, since the transition kernels $\rho_{\alpha}(x_{[\alpha]}^{\prime}|x_{[\alpha]};\epsilon)$ are not available in general, the weight vector $w_{\alpha}(x_{[\alpha]})\in\mathbb{R}^{d_{\alpha}}$ is approximated by the Schrödinger bridge approach as in (24). Algorithm 1 summarizes the localized Schrödinger bridge sampler.
Input: Input: Samples $\mathcal{X}\in\mathbf{R}^{d\times M}$.
Parameters : Bandwidth $\epsilon$. Localisation dimension $d_{\alpha}$. Desired number of new samples $N$. Number of decorrelation steps $n_{c}$.
Output: Output: New samples $x_{\rm s}^{(j)}$ for $j=1,\ldots,N$.

### v3

@@ 2.2. Kernel denoising and Tweedie’s formula
While (17) works well for low-dimensional problems and sufficiently large sample sizes $M$, applications to medium- or high-dimensional problems have remained an open challenge since accurate approximations of the Schrödinger bridge problem require an exponentially increasing number of samples as the dimension, $d$, of the sample space $\mathbb{R}^{d}$ increases (cf. [35]). The curse of dimensionality applies equally to the KDE-based approximation (35) and a failure to generalize has been discussed recently in the context of score-generative models [15].
@@
The key observation of this paper is that the approximation of conditional expectations (4) via Schrödinger bridges does not necessarily require the full Markov chain (12) and that localization can be applied provided conditional independence can be established. This idea will be developed in the following section. Localization will subsequently be extended to kernel-based denoising in Subsection 3.4.
@@ 3. Localized Schrödinger bridge sampler
To introduce the main idea of localizing the Schrödinger bridge sampler developed in [10] we first consider an illustrative example of sampling from a multivariate Gaussian distribution. We will see that localization allows for a significant reduction of the number of samples required to achieve a certain accuracy. In particular, the number of samples required to achieve a certain accuracy does not depend on the intrinsic dimension of the samples but rather is determined by the conditional independence which typically leads to a sequence of much lower dimensional estimation problems.
@@ 3.1. Motivational example: Gaussian setting
Let $\Delta_{h}\in\mathbb{R}^{d\times d}$ denote the standard discrete Laplacian over a periodic domain $[0,L]$ of length $L>0$ with mesh-size $h=L/d$. We assume that the sampling distribution $\pi(x)$ is Gaussian with zero mean and covariance matrix
(39) $C=(I-\Delta_{h})^{-1}.$
Instead of the distribution $\pi(x)$, we are given $M$ samples $x^{(j)}\sim{\rm N}(0,C)$, $j=1,\ldots,M$, and denote their $\alpha$-th entry by $x_{\alpha}^{(j)}$ for $\alpha=1,\ldots,d$. The goal is to produce more samples from ${\rm N}(0,C)$ using the time-stepping scheme (19) without making explicit reference to the unknown covariance matrix $C$. This particular setting of a generative model can become arbitrarily challenging by either increasing $L$ for fixed mesh-size $h$ or by decreasing the mesh-size $h=L/d$ for fixed $L$.
@@
(40) $X(n+1)=X(n)-\epsilon(I-\Delta_{h})X(n)+\sqrt{2\epsilon}\,\Xi(n),\qquad\Xi(n)\sim{\rm N}(0,I).$
@@
and periodic extension of $X_{\alpha}$ for $\alpha=0$ and $\alpha=d+1$. We assume that the step size $\epsilon$ is chosen such that $w_{0}\geq 0$. The EM update (41) reveals that the conditional expectation value of $X_{\alpha}(n+1)$ only depends on the value of the neighboring grid points of $X(n)$ with weights $w_{0}$ and $w_{\pm 1}$; i.e.,
@@
with $d_{\alpha}=3$, to denote the set of neighboring grid points of $X_{\alpha}$.
To help the reader navigating the various indices and sub- and superscripts we summarize here our notation. Superscripts $(j)$ are reserved to denote samples $j=1,\ldots,M$ as well as components of vectors in $\mathbb{R}^{M}$. For example, the components of the probability vector $w\in\mathbb{R}^{M}$ are denoted by $w^{(j)}$. The Greek subscript $\alpha$ with $\alpha=1,\ldots,d$ is reserved to denote components of a vector $x$ in state space $\mathbb{R}^{d}$, i.e. $x_{\alpha}$ for $\alpha=1,\ldots,d$. Subscripts ${[\alpha]}$ are reserved to denote localization around a component $\alpha$; i.e., $x_{[\alpha]}\in\mathbb{R}^{d_{\alpha}}$.
@@
for $\alpha=1,\ldots,d$. We recall from our previous considerations that the conditional expectation value of $X_{\alpha}(n+1)$ should depend on $X_{{[\alpha]}}(n+1/2)$ only. Hence the question arises whether we can find appropriately localized probability vectors $w(x)$ for the Schrödinger bridge sampler (19). The following formal argument can be made. We restrict ${\rm N}(0,C)$ to $x_{[\alpha]}\in\mathbb{R}^{d_{\alpha}}$ and truncate the samples $x^{(j)}$, $j=1,\ldots,M$, accordingly to yield $x_{{[\alpha]}}^{(j)}$. The covariance matrix $C_{\rm r}\in\mathbb{R}^{d_{\alpha}\times d_{\alpha}}$ of the reduced random variables $X_{{[\alpha]}}\in\mathbb{R}^{d_{\alpha}}$ is simply the restriction of $C$ to the corresponding sub-space, which in this particular example is independent of $\alpha$. Furthermore, using the Schur complement, one finds
(46) $C_{\rm r}^{-1}=\left(\begin{array}[]{ccc}\ast&-\frac{1}{2h^{2}}&\ast\\ -\frac{1}{2h^{2}}&1+\frac{1}{h^{2}}&-\frac{1}{2h^{2}}\\ \ast&-\frac{1}{2h^{2}}&\ast\end{array}\right),$
where $\ast$ denotes entries which differ from the matrix which would be obtained by restricting $C^{-1}$ to the corresponding sub-space. The important point is that the central elements remain identical (cf. (42)) and that only those entries enter the approximation of the conditional expectation value (43).
@@
(47) $(T_{\alpha})_{jk}=\exp\left(-\frac{1}{4\epsilon}\|x_{[\alpha]}^{(j)}-x_{[\alpha]}^{(k)}\|^{2}\right),\qquad j,k=1,\ldots,M,$
for fixed $\alpha\in\{1,\ldots,d\}$. For each of these localized matrices $T_{\alpha}$ we employ the local Sinkhorn algorithm to obtain the Sinkhorn weights $v_{\alpha}\in\mathbb{R}^{M}$ for $\alpha=1,\ldots,d$, which render
(48) $P_{\alpha}=D(v_{\alpha})T_{\alpha}D(v_{\alpha})$
bistochastic (cf.(13)). The key point is that the Euclidean norm in $\mathbb{R}^{d}$, $d\gg 1$, is replaced by the Euclidean norm in $\mathbb{R}^{d_{\alpha}}$ with $d_{\alpha}=3$. Furthermore, in this particular example, the corresponding Schrödinger bridge approximately couples the restricted Gaussian distribution ${\rm N}(0,C_{r})$ with itself. Next, the single $M$-dimensional probability vector (15) is replaced by $d$ $M$-dimensional probability vectors
@@
(51) $w^{(j)}_{\alpha}(x^{(k)}_{[\alpha]})=(P_{\alpha})_{jk},\qquad j,k=1,\ldots,M,$
@@
We obtain the localized approximation
(52) $m_{\alpha}(x_{[\alpha]};\epsilon)=\mathcal{X}_{\alpha}\,w_{\alpha}(x_{[\alpha]}),\qquad\alpha=1,\ldots,d,$
@@
(53) $\mathcal{X}_{\alpha}=(x_{\alpha}^{(1)},\ldots,x_{\alpha}^{(M)})\in\mathbb{R}^{1\times M}.$
We also introduce the localized data matrix
@@
For constant diffusion $S(x;\epsilon)=2\epsilon I$ the localized variant of the iteration scheme (5) becomes
(55) $X_{\alpha}(n+1)=m_{\alpha}(X_{[\alpha]}(n);\epsilon)+\sqrt{2\epsilon}\,\Xi_{\alpha}(n),\qquad\alpha=1,\ldots,d,$
for $\Xi(n)\sim{\rm N}(0,I)$. Similarly, the split-step scheme (19) becomes
(56a) $X_{[\alpha]}(n+1/2)$ $=X_{[\alpha]}(n)+\sqrt{2\epsilon}\,\Xi_{[\alpha]}(n),$
(56b) $X_{\alpha}(n+1)$ $=m_{\alpha}(X_{[\alpha]}(n+1/2);\epsilon).$
In other words, we have replaced a single Schrödinger bridge update in $\mathbb{R}^{d}$ by $d$ Schrödinger bridge updates in $\mathbb{R}^{d_{\alpha}}$.
Remark 3.
In line with the discussion on transformer architectures from Remark 2, we wish to point to a connection to multi-head attention [30]. More specifically, our localization procedure has introduced $d$ heads each relying on $\mathcal{X}_{[\alpha]}$ as matrix of key vectors, $\mathcal{X}_{\alpha}$ as matrix of value vectors, and $X_{[\alpha]}(n+1/2)$ as query vector in order to produce an update in the scalar-valued entries $X_{\alpha}(n+1/2)$ for $\alpha=1,\ldots,d$.
We finally discuss a localized version of the data-aware covariance matrix (18). Given localized weights $w_{\alpha}(x_{[\alpha]})$ and localized data matrices $\mathcal{X}_{[\alpha]}$, we define the $d_{\alpha}\times d_{\alpha}$-dimensional covariance matrices
(57) $S_{\alpha}(x_{[\alpha]};\epsilon)=\mathcal{X}_{[\alpha]}D(w_{\alpha}(x_{[\alpha]})\mathcal{X}_{[\alpha]}^{\top}\,-\mathcal{X}_{[\alpha]}w_{\alpha}(x_{[\alpha]})\,w_{\alpha}(x_{[\alpha]})^{\top}\mathcal{X}_{[\alpha]}^{\top}$
for $\alpha=1,\ldots,d$. Given a sample $\Xi(n)\sim{\rm N}(0,I)$, one first computes the $d_{\alpha}$-dimensional vector $\sqrt{S_{\alpha}(x_{[\alpha]};\epsilon)}\,\Xi_{[\alpha]}(n)$ of which one picks its scalar entry corresponding to $x_{\alpha}$, which we denote by $\xi_{\alpha}(n)$. The localized variant of (5) then becomes
(58) $X_{\alpha}(n+1)=m_{\alpha}(X_{[\alpha]}(n);\epsilon)+\xi_{\alpha}(n),\qquad\alpha=1,\ldots,d.$
Refer to caption       Refer to caption      Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 1. Comparison of the samples obtained from three different variants of localized Schrödinger bridge samplers. We show the centered rows of the empirical covariance matrix $\hat{C}$ (top row) and empirical histograms (bottom row) obtained from using all $d=101$ components. The blue markers denote the empirical covariance for the given samples; the magenta markers show the average over all $d$ rows. Left column: Localized EM-type sampler (55); middle column: Localized split-step sampler (56); right column: Localized sampler (58) with data-aware diffusion matrix (57). Given the large value of $\epsilon=1$, only (58) is able to faithfully reproduce the target measure ${\rm N}(0,C)$.
@@ 3.1.1. Numerical illustration
(59) $x^{(j)}\sim{\rm N}(0,C)$
for $j=1,\ldots,M$ with a $d\times d$ covariance matrix of the form (6) with a tridiagonal precision matrix with $C^{-1}_{i,i}=2$, $C^{-1}_{i,i\pm 1}=-0.5$, and periodic conditions $C^{-1}_{1,d}=C^{-1}_{d,1}=-0.5$. The corresponding entries of the covariance matrix are $C_{i,i}\approx 0.58$, $C_{i,i\pm 1}\approx 0.15$, $C_{i,i\pm 2}\approx 0.04$, and $C_{i,i\pm 3}\approx 0.01$.
We employ three different implementations of the localized Schrödinger bridge sampler with a localization set comprised of two neighboring grid points, i.e. $d_{\alpha}=3$. These implementations are (i) the split-step scheme (56), (ii) the localized EM-type scheme (55), and (iii) the scheme (58) with data-aware noise update.
In Figure 1 we compare the generated new samples with the given samples for all three sampling strategies. We show the resulting empirical histograms as well as the rows of the empirical covariance matrix $\hat{C}$. The rows are centered at the middle point using periodicity. We use $M=100$ training samples and a bandwidth of $\epsilon=1$ for $d=101$ to generate $10,000$ new samples.
While the split-step scheme (56) underestimates the variance of the distribution, the EM-type scheme (55) overestimates the variance. Only the localized scheme with data-aware diffusion (58) captures the marginal distribution and the covariance structure with desirable accuracy.
In Figure 2 we investigate the behavior of the localized split-step scheme (56) for varying parameter $\epsilon\in\{0.01,0.1,1\}$. We find that $\epsilon=0.1$ improves the results while a further decrease to $\epsilon=0.01$ leads to results comparable to $\epsilon=1.0$. Similar results are obtained for the EM-type sampling scheme with constant diffusion (55). An optimal choice of $\epsilon$ depends, of course, on the number of available training samples, which has been set to $M=100$ for these experiments.
We remark that unlocalized Schrödinger bridge samplers generate samples with a rather noisy correlation structure which is to be expected for $M=100$ training samples.
       Refer to caption       Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2. Comparison of the samples obtained from the localized split-step sampler (56) for varying parameter $\epsilon$. Left column: $\epsilon=1.0$; middle column: $\epsilon=0.1$; right column: $\epsilon=0.01$. While $\epsilon=0.1$ leads to improved results, it is found that larger and smaller values of $\epsilon$ degrade the performance of the split-step sampler.
@@ 3.2. Localized Schrödinger bridge sampler for general measures
The strategy of constructing a dimension-reduced localized Schrödinger bridge sampler as presented in the previous example of a multivariate Gaussian readily extends to general target measures $\nu({\rm d}x)$.
@@
The key assumption we make is that of conditional independence of $x_{\alpha}$ on $x_{\backslash{[\alpha]}}$, which allows for the dimension reduction. We consider variables $x_{\alpha}^{\prime}$ to be conditionally independent of $x_{\backslash{[\alpha]}}$ if the conditional distribution in $x_{\alpha}^{\prime}$, $p_{\alpha}(x_{\alpha}|x;\epsilon)$, satisfies
(60) $p_{\alpha}(x_{\alpha}^{\prime}|x;\epsilon)=p_{\alpha}(x_{\alpha}^{\prime}|x_{[\alpha]};\epsilon),$
which implies $\mathbb{E}[x_{\alpha}^{\prime}|x]=\mathbb{E}[x_{\alpha}^{\prime}|x_{[\alpha]}]$. The conditional expectation value (52) turns out to be a Monte-Carlo approximation of the conditional expectation under this assumption. More precisely, given the transition density of overdamped Langevin dynamics, denoted here by $p(x^{\prime}|x;\epsilon)$, we obtain
@@
where the second line follows from the conditional independence assumption. It is reasonable to assume that we can construct a reversible overdamped Langevin process with invariant distribution $\pi_{\alpha}(x_{[\alpha]})$ and transition kernel $p_{\alpha}(x_{[\alpha]}^{\prime}|x_{[\alpha]};\epsilon)$ on $\mathbb{R}^{d_{\alpha}}$. Here $\pi_{\alpha}(x_{[\alpha]})$ denotes the marginal distribution of $\pi(x)$ in $x_{[\alpha]}$. Then detailed balance of this dimension-reduced Langevin process is given by
@@
Note that in our localized Schrödinger bridge sampler detailed balance is ensured by the Sinkhorn algorithm which renders the Markov chain reversible. Detailed balance then implies
@@
We note that $\rho_{\alpha}(x_{[\alpha]}^{\prime}|x_{[\alpha]};\epsilon)$ is a density with respect to the reference measure induced by $\pi_{\alpha}(x_{[\alpha]})$. We finally approximate the integral in (63c) via Monte Carlo approximation using the restricted data samples $x_{[\alpha]}^{(j)}\sim\pi_{\alpha}(x_{[\alpha]})$, $j=1,\ldots,M$; i.e.,
@@
with $w_{\alpha}^{(j)}(x_{[\alpha]})\propto\rho_{\alpha}(x_{[\alpha]}^{(j)}|x_{[\alpha]};\epsilon)$ such that $\sum_{j}w_{\alpha}^{(j)}(x)=1$. This expression is of the form used in the localized Schrödinger bridge sampler (52). Furthermore, since the transition kernels $\rho_{\alpha}(x_{[\alpha]}^{\prime}|x_{[\alpha]};\epsilon)$ are not available in general, the weight vector $w_{\alpha}(x_{[\alpha]})\in\mathbb{R}^{d_{\alpha}}$ is approximated by the Schrödinger bridge approach as in (49). Algorithm 1 summarizes the localized Schrödinger bridge split-step sampler (56). The algorithm naturally extends to the sampler (58) with localized data-aware covariance matrix (57).
Input: Samples $\mathcal{X}\in\mathbf{R}^{d\times M}$.
Parameters : Bandwidth $\epsilon$. Localization dimension $d_{\alpha}$. Desired number of new samples $N$. Number of decorrelation steps $n_{c}$.
Output: New samples $x_{\rm s}^{(j)}$ for $j=1,\ldots,N$.